OpenStudy (anonymous):

if vector A=5i, B sub z=0, absolute value of vector B=7, vector A x vector B=2.5k, find vector B

OpenStudy (anonymous):

\[ \mathbf A=5\mathbf{i}+0\mathbf{j}+0\mathbf k \]

OpenStudy (anonymous):

\[ \mathbf B=b_x\mathbf{i}+b_y\mathbf{j}+0\mathbf k \]

OpenStudy (anonymous):

First do the cross product.

OpenStudy (anonymous):

but what pouts do i use for bi and bj

OpenStudy (anonymous):

Keep them as variables \(B_x\) and \(B_y\)

OpenStudy (anonymous):

i got o,o,5By

OpenStudy (anonymous):

Notice how: \[ 5B_y=2.5 \]

OpenStudy (anonymous):

Now you can find \(B_y\)

OpenStudy (anonymous):

you lost me, the 2.5 comes from b=k right?

OpenStudy (anonymous):

vector A x vector B=2.5k

OpenStudy (anonymous):

\[ \mathbf A\times \mathbf B =2.5\mathbf k \]

OpenStudy (anonymous):

\[ 2.5\mathbf k = 5B_y\mathbf k \]

OpenStudy (anonymous):

ok that was given...

OpenStudy (anonymous):

so By=.5?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Now: \[ \sqrt{B_x^2+B_y^2+0^2}=7 \]

OpenStudy (anonymous):

Because \[ \|\mathbf B\|=7 \]

OpenStudy (anonymous):

So \[ B_x^2+B_y^2=7^2 \]

OpenStudy (anonymous):

this is to find Bx?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

which I'm getting = 48.25

OpenStudy (anonymous):

Bx sqaured+.5^2=49 Bx=48.75?

OpenStudy (anonymous):

You need to do the square root.

OpenStudy (anonymous):

so im getting final answer of 6.98212i+.5j correct?

OpenStudy (anonymous):

thank you, i learned a different approach

OpenStudy (anonymous):

\[ |B_x|=\frac{\sqrt{195}}{2} \]

OpenStudy (anonymous):

There are two answers...

OpenStudy (anonymous):

\[ B_x=\pm \frac{\sqrt{195}}{2} \]

OpenStudy (anonymous):

i see, from the square root

OpenStudy (anonymous):

A different approach? There is only one approach I can think of. anyway we have the answer. there are two.

OpenStudy (anonymous):

I'm fairly new at this material

OpenStudy (anonymous):

Do you understand the work we did? Do you understand why there are two answers?

OpenStudy (anonymous):

yes, 2 answers because anytime you square root, you have plus answer and minus answer, and the work we did was, the way i see it, filling in the gaps from the info we got to complete to formula to find the answer